# Fishers Harvest Parallel Unlearning in Inherited Model Networks

### v1

@@ 4.3. Discussion
From the perspective of the FIM, the impact of data on specific model parameters is consistently localized at precise positions within the parameter space (Foster et al., 2024b). The stable localization of the FIM across iterations helps precisely locate the parameter locations for unlearning at the class, client, and sample levels using data tied to specific label sets. By fine-tuning $\tau$, $\eta$, and $\gamma$ in Eq. (6) to match the sample size of the class being unlearned, the task transits flexibly between the class-level (Foster et al., 2024b) and sample-level unlearning (Golatkar et al., 2020). Specifically, when the unlearning FIM mirrors the model’s FIM with a class-level setting, it becomes a client-level unlearning task. All types of unlearning are performed irrespective of the succession of inheritance relationships. Consequently, employing FIUn disrupts the sequential dependencies traditionally required for unlearning, thus enabling fully parallel unlearning processes across any number of subgraphs.
@@
Not only can models across different subgraphs undergo parallel unlearning, but those within the same subgraph at various depths can also be unlearned in parallel, thanks to a feature referred to as Hyper-Distance in this paper. This Hyper-Distance is achieved without compromising the integrity of the inheritance relationships. Hyper-Distance ensures that both unlearned and retained knowledge is effectively managed along the inheritance paths. This efficiency is possible because the FIUn method utilizes FIMs for unlearning, enabling a class-wise inheritance approach. Each model effectively incorporates a union of label sets from its predecessors, ensuring that the inheritance is preserved. Simultaneously, by focusing on label differences, we achieve fully parallel unlearning—making depths and paths irrelevant to the process.
@@
Answer to RQ2: The FIUn method enables independently removing knowledge from each inherited model. This method utilizes the FIM to obtain the Hyper-Distance property which breaks sequential dependencies among models, enabling fully parallel unlearning on inherited models and significantly decreasing the execution time of unlearning tasks. To address the complexity arising from multiple upstream models, the merging FIM function is developed to aggregate unlearning FIMs from those models, facilitating the efficient, one-shot removal of inherited knowledge.
@@
We also map the FIUn method and merging FIM function to the UMIG-assisted unlearning scenarios described in Section 3.3. This demonstrates how the proposed approach efficiently facilitates unlearning through the property of Hyper-Distance. Converting Imbalance into Balance. This scenario is particularly significant due to the fact that existing unlearning methods suffer from inevitable sequential dependencies. FIUn leverages the Hyper-Distance property to circumvent the influences of depths and paths within the inheritance topology. Consequently, imbalanced-path scenarios can be managed as effectively as balanced-path scenarios, where only one update operation is needed, regardless of whether unlearning is conducted synchronously or asynchronously, i.e., only one unlearning FIM needs to be computed for each inherited node.
@@
Converting Multi-root into Single-root. The merging FIM function is proposed to refine multi-root scenarios. This function merges the FIMs of the corresponding unlearned label sets from multiple discovery nodes into a single FIM, as shown in Figure 4. Specifically, overlapping inherited nodes require only one update operation to eliminate the impact of multiple discovery nodes by conducting a parameter-wise merger, where only the most impactful parameters take effect in this FIM. This function eliminates the need for multiple comparisons among various unlearning FIMs in multi-root scenarios, hence enhancing unlearning efficiency.
4.4. Implementation
4.4.1. System Setting
In practice, FIUn can be handled differently depending on the system architecture, as described in the following:
@@ 4.4.1. System Setting
  • Centralized: Model FIMs are computed and attached along with the model being published and sent to the central coordinator. When an unlearning request is initiated, the central server performs unlearning by merging the unlearning FIMs of the discovery nodes and executes unlearning in parallel.
@@
  • Decentralized: The discovery nodes publicly share Each unlearning FIM. Owners of each inherited node can then combine these unlearning FIMs with their own local models at will, according to specific unlearning requirements. This decentralized approach enables efficient and effective unlearning.
4.4.2. Complexity
@@
The computational complexity of an unlearning task comprises the complexities of Steps 1, 3, and 4 (i.e., calculate unlearning FIM, merge unlearning FIMs and update models). The complexity of Step 2 is not considered, as the computation of the model FIM can be performed beforehand. The overall complexity for the unlearning task is $\mathcal{O}(|\hat{D}||L|+|\{n^{\scriptscriptstyle D}_{i}\}|+|L|)$. Breaking down the complexities, Eq. (3) and Eq. (4) update $|L|$ parameters per data entry, with complexities $\mathcal{O}(|\hat{D}^{\scriptscriptstyle D}_{i}||L|)$ and $\mathcal{O}(|D_{j}||L|)$, respectively, where $|\hat{D}^{\scriptscriptstyle D}_{i}|\leq|\hat{D}|$ and $|L|$ is the number of parameters in the last layer. The term $\mathcal{O}(|D_{j}||L|)$ is omitted as model FIMs can be calculated along with the model training prior to any unlearning tasks. Eq. (5) includes a comparison and merger of $K$ unlearning FIMs, leading to a complexity of $\mathcal{O}(K)$, where $K$ is the number of traceable discovery nodes in the unlearning graph $\mathcal{\hat{G}}$ from node $n_{j}$ and $K\leq|\{n^{\scriptscriptstyle D}_{i}\}|$. Eq. (6) updates $|L|$ model parameters, leading to a complexity of $\mathcal{O}(|L|)$. As the computation of the unlearning FIM for discovery nodes and the model updates can proceed in parallel, the overall complexity of the unlearning tasks is $\mathcal{O}(|\hat{D}||L|+|\{n^{\scriptscriptstyle D}_{i}\}|+|L|)$.
5. Experimental Settings
@@
Table 1. Baseline of Federated Unlearning Performance On CIFAR-100 Unlearning Speed Comparison
Model #$C_{f}$ Cumulative Unlearning Time (s)
Gradient ascent Finetune Distill Re-training FIUn
@@
(a) FL in ResNet18 of #$C_{f}=1$
(b) FL in ResNet18 of #$C_{f}=10$
(c) FL in AlexNet of #$C_{f}=1$
(d) FL in AlexNet of #$C_{f}=10$
Figure 5. Baseline of Federated Unlearning Performance On CIFAR-100.
5.1. Datasets and Models
Datasets. The experiment uses two datasets to evaluate the proposed FIUn method. These two datasets are benchmarks for image classification tasks, encompassing diverse image classification tasks.
  • CIFAR100 (Krizhevsky et al., 2009). This dataset consists of 60,000 color images in 100 classes, with 600 images per class. The images are of size $32\times 32$ pixels. The dataset is divided into 50,000 training images and 10,000 test images. Each class in the CIFAR-100 dataset belongs to one of 20 superclasses. There is a coarse label (superclass) and a fine label (specific class) per image.
  • TinyImageNet (Le and Yang, 2015). This dataset contains 200 classes, each with 500 training images, 50 validation images, and 50 test images, with a total of up to 100,000 images. The images are resized to $64\times 64$ pixels, making them smaller than those in the original ImageNet dataset but larger than those in CIFAR-100.
Models. We use the AlexNet model (Iandola et al., 2016), ResNet18 model (Odusami et al., 2021) and DenseNet161 model (Talo, 2019) to evaluate the impact of our proposed FIUn method on the accuracy of the unlearned label set $C_{f}$ and the retained label set $C_{r}$.
Performance Metrics. We use the training datasets to check the accuracy of the model to evaluate its practicality in experiments. This is a common and reasonable approach since it directly evaluates if the removed information still influences the model, as widely accepted in the field (Tarun et al., 2023; Kim et al., 2024; Golatkar et al., 2020).
  • Accuracy on unlearned labels ($AD_{f}$). The accuracy of the unlearned label set in the unlearning model ideally would be close to zero. We use the training datasets to test the effectiveness of unlearning.
  • Accuracy on retained labels ($AD_{r}$). The accuracy of a retained label set of the unlearning model. It is expected to be close to the accuracy of the original model before unlearning.
  • Accuracy difference index ($\bigtriangleup_{acc}$). The difference between $AD_{r}$ and $AD_{f}$ also serves as a direct metric to evaluate unlearning performance. The greater the difference, the more effective the unlearning is, with a maximum 1.
Table 2. Federated Unlearning Performance On TinyImageNet
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
@@ 5.1. Datasets and Models
DenseNet161 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 88.03 77.64 76.02 999.41 1996.43 2997.64 2.51 4.59 4.95
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.01 71.43 71.00 1004.64 2006.31 3006.14 3.16 4.69 4.89
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 80.00 67.67 66.58 1019.43 2020.16 3018.64 3.10 4.96 4.75
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.72 99.71 99.99 99.99 71.68 58.01 55.80 1031.43 2029.64 3030.64 2.95 4.85 4.89
${AD}_{f}\downarrow$ 99.99 99.99 98.82 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.73 99.97 99.46 112.95 216.39 314.29 1.35 3.68 3.42
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.17 99.98 99.98 109.89 209.31 310.42 1.63 3.26 3.57
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 99.54 99.98 99.98 107.39 208.12 305.32 1.46 3.37 3.51
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.98 99.99 99.99 104.21 201.36 302.74 1.52 3.21 3.36
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
Table 3. Incremental Unlearning Performance On TinyImageNet
@@
Model #$C_{f}$ Metrics Original Model (%) Re-training Model (%) FIUn (%) Cumulative Unlearning Time (s)
@@
DenseNet161 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.81 99.99 99.99 98.80 89.76 78.50 73.87 991.43 1995.36 3989.47 2.75 4.85 4.43
${AD}_{f}\downarrow$ 99.99 99.99 99.66 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.80 99.99 99.99 99.51 83.98 65.87 59.11 1006.32 2004.75 4009.41 2.64 4.43 4.64
${AD}_{f}\downarrow$ 99.99 99.99 99.49 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.96 99.99 99.99 99.16 86.81 61.02 55.04 1019.43 2019.50 4019.46 2.36 4.47 4.74
${AD}_{f}\downarrow$ 99.99 99.99 99.89 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.72 99.99 99.99 99.99 85.95 54.61 45.46 1035.36 2041.78 4039.75 2.65 4.46 4.75
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
ResNet18 1 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.66 96.33 96.07 120.32 221.75 439.85 1.37 3.75 3.43
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
4 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 98.84 96.34 96.70 115.32 217.33 435.17 1.41 3.34 3.64
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
6 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 98.82 99.99 98.59 110.32 211.35 431.37 1.47 3.46 3.75
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 ${AD}_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.33 97.51 98.45 108.32 207.43 428.43 1.32 3.56 3.57
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
5.2. Learning Framework Setup and Benchmarks
Figure 6. Models topologies considered in the experiments.
Targeted Frameworks. We detailed setups for four learning frameworks under the UMIG with varying dataset and label distribution conditions.
  • Federated Unlearning. This process uses the structure shown in Figure 6. The CIFAR100 and TinyImageNet datasets are each randomly divided into five parts. When the model $w_{g}$ contains labels that need to be unlearned, the models that inherit from it need to perform the unlearning operation. We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
  • Distributed Data-Parallel Unlearning. This process uses the structure shown in Figure 6. The CIFAR100 and TinyImageNet datasets are each randomly divided into three parts, each executing a sub-task in parallel. When the model $w_{g}$ contains labels that need to be unlearned, each sub-task performs the unlearning operation in parallel. We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
  • Incremental Unlearning. This process uses the structure shown in Figure 6. In the CIFAR100 dataset, the first model is trained on data with labels 0-90, and each inheriting model’s training dataset progressively includes two additional classes. In the TinyImageNet dataset, the first model is trained on data with labels 0-190, and similarly, each inheriting model’s training dataset progressively includes two additional classes. We select models $w_{g}$, $w_{a}$, and $w_{b}$ for experimental demonstration.
  • Transfer Unlearning. Since TL involves fine-tuning the last layer parameters, adjustments are focused on the last layer of the model. Experiments, as illustrated in Figure 6, involve the CIFAR100 dataset where a model $w_{g}$ is trained on labels 0-90 and then transferred to models $w_{a}$ and $w_{b}$ trained on labels 0-92 and 0-98, respectively. Similarly, for the TinyImageNet dataset, model $w_{g}$ is trained on labels 0-190 and then transferred to models $w_{a}$ and $w_{b}$, which are trained on labels 0-192 and 0-198, respectively.
@@ 5.2. Learning Framework Setup and Benchmarks
Benchmark. The experiment specifically focuses on class-level unlearning tasks for clarity of presentation. We re-train the models that include unlearned labels and subsequent models influenced by these labels due to inheritance. We remove the data with unlearned labels from each model’s dataset. The remaining data is then used to individually train each model. The models are trained sequentially according to their inheritance relationships.
Hyperparameters. The hyperparameters used in the experiment are shown in Appendix .1. The table lists the learning rates, optimizers, and other hyperparameters for different models, as well as the hyperparameters of our proposed FIUn method. Throughout the experimental process, all the hyperparameters remain unchanged to ensure a fair comparison of our method among different models.
Table 4. Transfer Unlearning Performance On CIFAR-100

### v4

@@ V-B Discussion
From the perspective of the FIM, the impact of data on specific model parameters is consistently localized at precise positions within the parameter space [foster2024fast]. The stable localization of the FIM across iterations helps precisely locate the parameter locations for unlearning at the class, client, and sample levels using data tied to specific label sets. By fine-tuning $\tau$, $\eta$, and $\gamma$ in (6) to match the sample size of the class being unlearned, the task transits flexibly between the class-level [foster2024fast] and sample-level unlearning [golatkar2020eternal]. Specifically, when the unlearning FIM mirrors the model’s FIM with a class-level setting, it becomes a client-level unlearning task. All types of unlearning are performed irrespective of the succession of inheritance relationships. Consequently, employing FIUn disrupts the sequential dependencies traditionally required for unlearning, thus enabling fully parallel unlearning processes across any number of subgraphs.
@@
Models across subgraphs and at different depths within the same subgraph can undergo parallel unlearning, enabled by Hyper-Distance, where learned knowledge’s parameter impact is independent of model inheritance depth. Each model integrates predecessors’ label sets to preserve inheritance; focusing on label differences achieves full parallel unlearning, making depths/paths irrelevant without compromising inheritance integrity. This is possible because the FIUn method utilizes FIMs for unlearning, enabling a class-wise inheritance approach.
@@
Answer to RQ2: The FIUn method enables the independent removal of knowledge from each inherited model. This method utilizes the FIM to obtain the Hyper-Distance property which breaks sequential dependencies among models, enabling fully parallel unlearning on inherited models and significantly decreasing the execution time of unlearning tasks. To address the complexity arising from multiple upstream models, the MFIM function is developed to aggregate unlearning FIMs from those models, facilitating the efficient, one-shot removal of inherited knowledge.
We also map the FIUn method and the MFIM function to the DAG-assisted unlearning scenarios described in Sec. IV-C. This shows how the proposed method efficiently facilitates unlearning through the property of Hyper-Distance.
@@
Converting Imbalance into Balance. This scenario is particularly significant since existing unlearning methods face unavoidable sequential dependencies. FIUn utilizes the Hyper-Distance property to bypass the effects of depths and paths in the inheritance topology. As a result, imbalanced-path scenarios are handled as efficiently as balanced ones, requiring just one update operation per inherited node, regardless of whether unlearning is synchronous or asynchronous.
@@
Converting Multi-root into Single-root. The MFIM function is proposed to refine multi-root scenarios. This function merges the FIMs of the corresponding unlearned label sets from multiple discovery nodes into a single FIM, as shown in Fig. 4. Specifically, overlapping inherited nodes require only one update operation to eliminate the impact of multiple discovery nodes by conducting a parameter-wise merger, where only the most impactful parameters take effect in this FIM. This function eliminates the need for multiple comparisons among various unlearning FIMs in multi-root scenarios, hence enhancing unlearning efficiency.
V-C Implementation
System Settings. FIUn implementation varies based on system architecture:
  • $\bullet$
@@ V-C Implementation
    Centralized: Model FIMs are computed and sent with the published model to the central server. Upon an unlearning request, the central server merges the unlearning FIMs from the discovery nodes and executes parallel unlearning.
  • $\bullet$
@@
    Decentralized: Each unlearning FIM is publicly shared by discovery nodes. Inherited node owners can then combine these FIMs with their local models as needed, enabling efficient and effective decentralized unlearning.
Although FIUn focuses on efficient parallel unlearning, it can be extended with auditability mechanisms to ensure verifiable unlearning in both centralized and decentralized settings. A practical approach is to integrate FIM commitments that record the Fisher information signatures before and after unlearning. Each node can compute a lightweight hash or zero-knowledge proof of its local FIM difference $\Delta F=F-\hat{F}^{D}$, which is then submitted to a verification server or blockchain ledger. In a centralized mode, the server validates updates by comparing FIM commitments to the expected parameter deltas from (6). In a decentralized mode, commitments are written on-chain, enabling independent validators to check that unlearning matches the declared forgetting requests without revealing model parameters. This yields an auditable trail for the “right to be forgotten” while preserving FIUn’s efficiency and scalability.
Complexity. The computational complexity of an unlearning task comprises the complexities of Steps-1, 3, and 4 (i.e., calculate unlearning FIM, merge unlearning FIMs and update models, cf. Section V-A). The complexity of Step-2 is not considered, as the computation of the model FIM can be performed beforehand.
[…]
@@
Equation (5) includes a comparison and merger of $K$ unlearning FIMs, leading to a complexity of $\mathcal{O}(K)$, where $K$ is the number of traceable discovery nodes in the unlearning graph $\mathcal{\hat{G}}$ from node $n_{j}$ and $K\leq|\{n^{\scriptscriptstyle D}_{i}\}|$. Equation (6) updates $|L|$ model parameters, leading to a complexity of $\mathcal{O}(|L|)$. As the computation of the unlearning FIM for discovery nodes (including the forward propagation) and the model updates can proceed in parallel, the overall complexity of the unlearning tasks is $\mathcal{O}(|\hat{D}||L_{\mathrm{all}}|+|\hat{D}||L|+|\{n^{\scriptscriptstyle D}_{i}\}|+|L|)$.
VI Experimental Settings
@@
Figure 5: Models topologies considered in the experiments.
TABLE I: Hyperparameters.
Defaults (all models unless overridden): optimizer = SGD, learning rate = 0.01, epochs = 100, batch size = 64.
Overrides:
FIUn: lr = 0.1; $\tau$ = 1; $\gamma$ = 1; $\eta$ = 0.1; use layer = last.
AlexNet: epochs = 200.
BERT: optimizer = Adam; lr = 2e-5; epochs = 30.
GCNN: optimizer = Adam; lr = 0.001; epochs = 30.
ResNet18: lr = 0.1.
TABLE II: Baseline of Various Framework Performance on CIFAR100 Unlearning Speed Comparison
Model
Framework
$C_{f}$ Cumulative Unlearning Time (s)
Gradient Ascent Fine-tuning Distill Re-training FIUn
@@
IL 1 1.52 1.63 1.84 1.35 2.65 3.86 98.51 170.66 317.53 26.35 47.12 66.93 0.30 0.80 0.85
10 1.33 1.45 1.62 1.47 2.24 3.74 81.74 143.53 275.34 18.93 39.93 58.93 0.32 0.88 0.89
DDPL 1 1.18 1.25 1.34 1.57 1.43 1.51 72.34 71.87 71.39 32.53 31.95 32.18 0.64 0.34 0.30
10 1.34 1.42 1.51 2.04 1.85 1.97 71.41 66.56 70.41 22.60 21.92 24.19 0.68 0.34 0.34
TL 1 1.39 1.48 1.52 1.37 3.67 3.25 95.4 163.54 168.51 22.21 41.52 42.93 0.15 0.24 0.24
10 1.35 1.51 1.46 1.5 5.53 5.54 81.15 145.73 151.52 19.10 42.60 41.36 0.16 0.27 0.27
TABLE III: Transfer Unlearning Performance on CIFAR100
Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
@@
AlexNet
1 $AD_{r}\uparrow$ 97.16 95.61 92.20 98.17 95.98 91.32 91.40 86.24 83.99 20.31 40.49 42.36
${AD}_{f}\downarrow$ 96.66 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.09 0.11 0.11
2 $AD_{r}\uparrow$ 97.16 95.61 92.20 96.88 96.68 93.16 82.36 80.47 79.99 21.20 40.56 41.92
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.08 0.13 0.11
4 $AD_{r}\uparrow$ 97.16 95.61 92.20 94.67 96.45 93.45 71.59 71.43 76.06 20.30 39.76 40.60
${AD}_{f}\downarrow$ 96.61 98.30 98.30 0.00 0.00 0.00 0.00 0.00 0.00 0.09 0.11 0.11
ResNet18
1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.09 93.51 95.22 22.21 41.52 42.93
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.15 0.24 0.24
2 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 85.89 94.89 95.79 21.19 41.72 43.20
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.17 0.23 0.26
4 $AD_{r}\uparrow$ 99.96 99.99 99.99 99.99 99.99 99.99 80.20 94.82 97.03 22.03 42.26 43.25
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.17 0.25 0.25
VI-A Datasets and Models
Datasets. The experiment uses three datasets to evaluate the proposed FIUn method. These three datasets are benchmarks for image classification tasks, encompassing diverse image classification tasks.
  • $\bullet$
    CIFAR100. This dataset includes 60,000 color images across 100 classes, each with 600 images, sized at $32\times 32$ pixels. It comprises 50,000 training and 10,000 test images. Each class belongs to one of 20 superclasses, with both a coarse (superclass) and fine (specific class) label per image.
  • $\bullet$
    TinyImageNet. This dataset includes 200 classes, with 500 training images, 50 validation images, and 50 test images per class, totaling 100,000 images. The images are resized to $64\times 64$ pixels, larger than CIFAR100 but smaller than the original ImageNet.
  • $\bullet$
    Yahoo! Answers. his text dataset is a large public Q&A dataset widely used in natural language processing (NLP) research for tasks like text classification, question answering, and sentiment analysis. It contains categorized question-answer pairs from Yahoo! Answers.
Models. We use the AlexNet, ResNet18, DenseNet161, BERT-base-cased language model and GCNN to evaluate the impact of our proposed FIUn method on the accuracy of the unlearned label set $C_{f}$ and the retained label set $C_{r}$.
Performance Metrics. We use the training datasets to check the accuracy of the model to evaluate its practicality in experiments. This is a common and reasonable approach since it directly evaluates if the removed information still influences the model, as widely accepted in the field [tarun2023fast, golatkar2020eternal].
  • $\bullet$
    Accuracy on unlearned labels ($AD_{f}$). The accuracy of the unlearned label set in the unlearning model should ideally approach 0%. This represents a scenario where the softmax layer assigns negligible probabilities to unlearned classes. When applying the Top-$N$ strategy, the likelihood of selecting unlearned classes becomes extremely small, effectively resulting in 0% accuracy. We evaluate this by testing the unlearning model on training datasets to assess the effectiveness of the unlearning process.
  • $\bullet$
    Accuracy on retained labels ($AD_{r}$). The accuracy of the retained label set should closely match the original model’s accuracy before unlearning.
  • $\bullet$
  • $\bullet$
    Accuracy difference index ($\bigtriangleup_{acc}$). The difference between $AD_{r}$ and $AD_{f}$ directly measures unlearning effectiveness, with a larger difference indicating more effective unlearning, up to a maximum value of 1.
(a) FL in ResNet18 of $C_{f}=1$
(b) FL in ResNet18 of $C_{f}=10$
(c) FL in AlexNet of $C_{f}=1$
(d) FL in AlexNet of $C_{f}=10$
(e) DDPL in ResNet18 of $C_{f}=1$
(f) DDPL in ResNet18 of $C_{f}=10$
(g) DDPL in AlexNet of $C_{f}=1$
(h) DDPL in AlexNet of $C_{f}=10$
Figure 6: Baseline Unlearning Accuracy of FL and DDPL Frameworks on CIFAR100.
(a) IL in ResNet18 of $C_{f}=1$
(b) IL in ResNet18 of $C_{f}=10$
(c) IL in AlexNet of $C_{f}=1$
(d) IL in AlexNet of $C_{f}=10$
(e) TL in ResNet18 of $C_{f}=1$
(f) TL in ResNet18 of $C_{f}=10$
(g) TL in AlexNet of $C_{f}=1$
(h) TL in AlexNet of $C_{f}=10$
Figure 7: Baseline Unlearning Accuracy of IL and TL Frameworks on CIFAR100.
VI-B Framework Setup and Benchmarks
Targeted Frameworks. We detail setups for four learning frameworks with DAG, all adopting the structure shown in Fig. 5. These frameworks vary in their dataset and label distribution conditions.
  • $\bullet$
    Federated Unlearning. (Yellow in Fig. 5) The CIFAR100 and TinyImageNet datasets are each randomly divided into five parts. When the model $w_{g}$ contains labels that need to be unlearned, the models $w_{a}$ and $w_{b}$ that inherit from it need to perform the unlearning operation.
  • $\bullet$
    Distributed Data-Parallel Unlearning. (Red in Fig. 5) The CIFAR100 and TinyImageNet datasets are each randomly divided into three parts, each executing a sub-task in parallel. When the model $w_{g}$ contains labels that need to be unlearned, sub-tasks $w_{a}$ and $w_{b}$ perform the unlearning operation in parallel.
  • $\bullet$
    Incremental Unlearning. (Blue in Fig. 5) In the CIFAR100 and TinyImageNet datasets, the initial models are trained on labels 0-90 and 0-190, respectively, with each inheriting model adding two more classes. Models $w_{g}$, $w_{a}$, and $w_{b}$ are selected for experimental demonstration.
  • $\bullet$
    Transfer Unlearning. (Red in Fig. 5) Since TL involves fine-tuning the last layer, adjustments are focused on this layer. In experiments, a model $w_{g}$ is trained on CIFAR100 labels 0-90, then transferred to models $w_{a}$ and $w_{b}$ trained on labels 0-92 and 0-98, respectively. Similarly, in TinyImageNet, model $w_{g}$ is trained on labels 0-190 and transferred to models $w_{a}$ and $w_{b}$ trained on labels 0-192 and 0-198.
TABLE IV: Distributed Data-Parallel Unlearning Performance on CIFAR100
@@ VI-B Framework Setup and Benchmarks
Model $C_{f}$ Metrics Original (%) Re-training (%) FIUn (%) Cumulative Unlearning Time (s)
@@
AlexNet
1 $AD_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 97.23 98.85 97.66 30.13 31.26 31.14
${AD}_{f}\downarrow$ 99.99 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00 0.23 0.14 0.09
2 $AD_{r}\uparrow$ 99.96 99.96 99.96 99.96 99.96 99.96 96.14 97.07 96.71 26.43 26.47 26.41
${AD}_{f}\downarrow$ 99.96 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00 0.27 0.16 0.11
4 $AD_{r}\uparrow$ 98.01 99.96 99.96 99.96 99.96 99.96 94.58 90.41 90.97 24.33 24.36 24.37
${AD}_{f}\downarrow$ 99.09 99.96 99.96 0.00 0.00 0.00 0.00 0.00 0.00 0.23 0.14 0.10
ResNet18
1 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 91.87 94.67 93.67 32.53 31.95 32.18
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.64 0.34 0.30
2 $AD_{r}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 89.94 94.58 92.40 28.28 27.22 28.47
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.66 0.34 0.30
4 $AD_{r}\uparrow$ 99.96 89.20 88.52 99.96 99.96 86.44 93.95 89.26 91.77 25.37 24.97 25.13
${AD}_{f}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00 0.66 0.35 0.31
@@
Benchmark. The experiment specifically focuses on class-level unlearning tasks for clarity of presentation. We re-train the models that include unlearned labels and subsequent models influenced by these labels due to inheritance. We remove the data with unlearned labels from each model’s dataset. The remaining data is then used to train each model individually. The models are trained sequentially according to their inheritance relationships.
Hyperparameters. Table I summarizes the hyperparameters for various models and our FIUn method. All hyperparameters remain fixed during experiments to ensure fair comparisons across models.
VI-C Label Category Settings
The unlearned label categories, denoted as $C_{f}$, may vary across different models and data distributions. For instance, model $w_{a}$ aims to unlearn labels 2, 3, and 4, while model $w_{b}$ intends to unlearn labels 2, 3, and 6. There is an overlap between the unlearned label categories of model $w_{a}$ and model $w_{b}$, specifically labels 2 and 3. We refer to the degree of overlapping among these unlearned label categories as $\cap_{f}$. The unlearning process is applied to individual and combined labels. For CIFAR100 and Yahoo! Answers, we select combinations of 1, 2, and 4 categories, while for TinyImageNet, we select 1, 4, and 6 categories. The overlap of unlearned labels is divided into 60%, 40%, and 20%.
VII Evaluation
We use an NVIDIA 4090 GPU to evaluate FIUn’s impact on the accuracy of unlearned and retained labels across four frameworks: FL, DDPL, IL, and TL. Accuracy and time consumption during unlearning are assessed by comparing FIUn with existing methods, followed by an analysis of re-training, the most effective (but inefficient) unlearning benchmark, focusing on inheritance depth and parameter layers impact on unlearning effectiveness.
VII-A General Label Unlearning Analysis
We consider existing emerging unlearning methods for comparison: 1) fine-tuning, which adjusts the original model using the remaining dataset [odusami2021analysis], 2) gradient ascent (GA), employing negative gradients for unlearning [ga], and 3) distill, which distills knowledge from the original model into a student model using the remaining data [distil1Unl].
Table II reports unlearning speed and Fig.6 shows unlearning accuracy $\bigtriangleup_{acc}$; Fig.7 compares $\bigtriangleup_{acc}$ under IL and TL. Our method improves unlearning speed by up to 99% over alternatives and maintains a consistent accuracy advantage: for single-label and up to 10 classes, GA and fine-tuning finish in about 2 seconds, distillation and retraining are slower, while our approach averages under 1 second. In accuracy, fine-tuning and GA remain suboptimal, and our method matches or surpasses distillation across most architectures (with a slight dip on AlexNet). Overall, it delivers the fastest cumulative unlearning across IMNs. Building on this, we next compare with retraining, which is widely regarded as the gold standard for unlearning and retention accuracy [sisa]; see Tables III and VI, with additional results in Appendix A.
TABLE V: Federated Unlearning Performance on TinyImageNet